# Strategic Effort and Bandwagon Effects in Finite Multi-Stage Games with Non-Linear Externalities: Evidence from Triathlon using a Leave-One-Out IV Strategy

### v1

@@ Abstract
This paper examines strategic effort and positioning choices resulting in bandwagon effects under externalities in finite multi-stage games using causal evidence from triathlon (Reichel , 2025). Focusing on open-water swim drafting—where athletes reduce drag most effectively by swimming directly behind peers—we estimate its performance effects through a structural contest framework with endogenous, deterministic effort and drafting position. Leveraging exogenous variation from COVID-19 drafting bans in Austrian triathlons, we apply a panel leave-one-out (LOO/LOTO) peer ability instrumental variables (IV) strategy to isolate the causal non-linear effect of drafting. Results from restricted sample analysis and pooled estimated bandwagon IV effects show substantial and nonlinear gains: in small (group size $<$ 10) drafting swim groups/clusters, each deeper position improves finishing rank on average by over 30%, with rapidly diminishing returns in larger groups. Leading however is consistently more costly than optimal positioning, aligning with theoretical predictions of energy expenditure (metabolic costs).
@@ 1 Introduction
This paper investigates the performance implications of open water swim drafting using rich panel data from triathlon races by modeling drafting in the first segment as a two-stage deterministic game. In the first stage, athletes exert some deterministic swimming effort; in the second stage, they position themselves optimally within a group, potentially benefiting heterogeneously from reduced drag behind other swimmers due to the inversly decaying nature of hydrodynamic benefits and rare overtakings (resorting) during the swim segment (Reichel , 2025, 2025).
Empirical estimates of drafting effects in the swim segment are still limited by measurement challenges. Group formations are likely endogenous and cannot be observed directly, and drafting exposure may correlate with unobserved ability (sorting/selection bias). Laboratory and biomechanical studies (e.g., Chatard , 1990, 2003; Delextrat , 2003; bolon2020drafting) focus on energy expenditure (metabolic) and performance gains from drafting, but large-scale causal estimates from race data have only been documented recently (in Reichel , 2025).
In our theoretical framework, we extend contest theory and incorporate models of strategic externalities (Tullock , 1980; Moldovanu , 2001; Konrad , 2009; Che , 2003) to develop a structural model in which athletes first select a deterministic level of swimming effort, followed by an endogenous choice of drafting position. Drafting benefits are captured through a non-linear function that exhibits diminishing marginal returns beyond a certain positional threshold. The structural implications of the model are then integrated into an empirical two-stage least squares (2SLS) estimation framework. To address endogeneity in within-group drafting positions, we employ a leave-one-out (LOO/LOTO) average of peer swimming ability as an instrumental variable (IV).
@@ 2.2 Drafting Group Construction
Drafting groups can not be observed directly in the data. We thus infer drafting clusters using a single-linkage hierarchical agglomerative clustering algorithm approach applied to swim-out segment times within each event. Athletes are grouped into the same swim cluster if their exit times fall within a 5-second distance thresholds recorded. Earlier trials with complete linkage and wider thresholds (10–15 seconds) produced overly broad or unstable groupings. (Reichel , 2025)
@@ 3.1 A Two-Stage Finite Game
This structure reflects a dynamic contest framework with endogenous positioning and group-related externalities (see Tullock , 1980; Konrad , 2009; Moldovanu , 2001; Tadelis , 2013; Vives , 2005).
@@ 3.1.1 Effort Cost Function
For simplicity we assume a linear cost function of effort:
@@ 3.2.2 Second Derivative and Convexity
3.2.3 Proposition (Interior Optimal Drafting Position):
Proposition 1: There exists a unique $d_{i}^{*}>3$ such that disutility is minimized, i.e., drafting benefit is maximized net of performance shortfall risk.
Proof: - For $d_{i}\leq 3$, $B(d_{i})=0\Rightarrow$ disutility is flat. - For $d_{i}>3$, $DIS_{i}$ is strictly convex. - Since $DIS_{i}\to\alpha c\bar{e}_{i}^{1}-(1-\alpha)\gamma$ as $d_{i}\to\infty$, the marginal gains vanish. - Hence, the minimum is interior and finite.
@@ 3.2.4 Empirical Mapping
or to estimate effects. We can address address endogeneity by instrumenting $d_{i}$ using a leave-one-out peer ability IV.
@@ 3.3 Why the Optimal Position Is Not the Last
Conclusion: Athletes face trade-offs — drafting increases with depth, but so do tactical disadvantages. Therefore, the optimal position is interior — balancing maximum benefit with control. This matches observed behavior in collective movement and contests with crowding externalities (Hillman , 1989; Krueger , 1974).
@@ 4.1 Research Objective
Because drafting positions reflect strategic sorting and group formation dynamics, naive OLS is likely biased. We thus use a two-stage least squares (2SLS) framework with a leave-one-out peer swim-time instrument to estimate a local average treatment effect (LATE) of drafting benefits (functional form from the theoretic model) or drafting positions on performance.
@@ 4.3 Endogeneity and Identification Strategy
We address this using a constructed leave-one-out peer swim-time instrument $Z_{ie}$ to generate the relevant and exogenous variation in swim group drafting structure.
@@ 4.3.1 Leave-One-Out (LOO/LOTO) Instrumental Variable (IV)
We, define the instrument as:
@@
This captures the swim ability of peers in athlete $i$’s drafting within the current group, excluding $i$.This leave-one-out (LOO/LOTO) instrument captures the average peer ability while avoiding mechanical reflection bias. By excluding athlete i, we remove the mechanical correlation between the instrument and the outcome. This strategy aligns with methods used in social interaction and peer effect models to address endogeneity, e.g., Bramoulle (2009) and Angrist (2014). Related applications include De Loecker (2011), who employ LOO instruments in production function estimation, and Guryan (2008), who instrument for peer characteristics using group averages excluding the individual.
@@ 4.3.2 Drafting Benefit Transformation from Theoretic Model
4.4 Empirical Specification: Two-Stage Least Squares (2SLS) Peer-IV Estimation
To identify the causal effect of within-group drafting benefit on race performance, we estimate a two-stage least squares (2SLS) model using a leave-one-out peer ability instrument. This setup accounts for endogeneity in observed drafting positions, which are determined by both athlete ability and strategic group formation.
@@ 4.4 Empirical Specification: Two-Stage Least Squares (2SLS) Peer-IV Estimation
We begin by estimating the drafting position as a function of peer swim ability and controls:
@@
  • $Z_{ie}$: leave-one-out peer average swim time (excluding athlete $i$);
@@
The instrument $Z_{ie}$ introduces exogenous variation by exploiting group composition. Since athletes do not control who else is in their drafting cluster, this measure captures plausibly exogenous shifts in strategic positioning pressure.
@@
Using the fitted drafting benefit $\widehat{B}_{ieg}=f(\widehat{D}_{ieg})$ to validate the theoretical model or using $\widehat{D}_{ieg})$ directly, we estimate:
@@
Identification Strategy and Exogeneity
This 2SLS framework identifies the local average treatment effect (LATE) of drafting benefit for compliers—athletes whose position is affected by peer group composition. Identifying Assumptions are:
@@
  • Instrument exogeneity: $\mathbb{E}[\varepsilon_{ieg}\mid Z_{ie},S_{ie},\bm{X}_{ie},\lambda_{e},\alpha_{i},\alpha_{g}]=0$. – $Z_{ie}$;
  • Monotonicity: An increase in peer ability (faster peers) weakly shifts athletes to deeper drafting positions.
@@
where: - $\widehat{B}_{ieg}$: instrumented drafting benefit; - $leader_{ieg}$: indicator for leading position ($D=1$); - $\lambda_{e},\alpha_{i},\alpha_{g}$: event, athlete, and group fixed effects.
@@ 4.4.1 Structural Validation and Interpretation
The coefficient $\beta_{1}$ reflects the average marginal return to drafting benefit. The increasing and concave structure implies that: - Greatest positive net gains occur when transitioning from front positions backwards; Diminishing returns occur deeper into the pack; The optimal position is interior, as predicted by the theoretical model.
Empirical evidence of this shape supports the game’s interior equilibrium logic and nonlinear (dis-)utility (drafting befenefit structure) topology.
@@ 4.5 Potential Extension: Projected Panel IV
For robustness and within-athlete variation, we define a projected instrument using historical group characteristics from the largest group containing athlete ’i’ f.e.:
@@ 5 Empirical Results
In Table 3, the coefficient on drafting position is $-0.011$ ($p<0.1$) in the full sample (without restricting the outcome variable), implying that moving one position deeper in the drafting line improves expected rank by approximately $1.1\%$:
@@
This effect increases substantially in smaller groups. In drafting clusters with fewer than 10 athletes, the coefficient becomes $-0.972$ ($p<0.01$), translating to a $62.1\%$ improvement in finishing rank per position:
@@
In very small clusters (group size $<5$), the coefficient is $-2.378$ ($p<0.1$), corresponding to an $90.7\%$ improvement:
@@ 5.1 Using the theory-based functional form
with some uncalibrated $\lambda=0.5$ for the simple sake of simplicity and brevity. This specification reflects increasing but diminishing returns to deeper drafting positions. Across all models, the coefficient on $B(d_{i})$ is significantly negative, consistent with improved performance from stronger drafting effects. The effect is strongest in smaller groups: in groups with fewer than 10 athletes, a one-unit increase in drafting benefit improves log-rank by approximately 7.1 to 9.6 points (Columns 5–6). In broader samples, the benefit remains significant and negative—$-6.65$ in the full sample (Column 1), and $-4.84$ to $-3.61$ in groups smaller than 20 (Columns 3–4).
@@ 6 Conclusion
This paper contributes comprehensive analysis on Strategic Effort and Bandwagon Effects in Finite Multi-Stage Games with Non-linear Externalities and draws on empirical causal evidence using rich panel data from drafting in endurance sports competitions, predicting effects on final stage rank outcomes focusing on competitive athletes. We develop and estimate log-linear models with endogenous drafting position, using a leave-one-out peer swim ability as an instrument (Angrist , 2009; Manski , 1993; De Giorgi , 2020). The results validate the strategic and nonlinear nature of drafting benefits in multi-stage competitions such as triathlon.
Our key findings are:
@@ Appendix A Appendix
A.1 Leave-one-out Peer-IV 2SLS Regressions with Capped/Limited Outcome Variable Rank $<$ 250
@@
(1) Full Sample (2) Group Size $<$ 10 (3) Group Size $<$ 5 Drafter Position (Fit) -0.011 -0.972∗∗∗ -2.378 (0.0053) (0.185) (1.073) Observations 34,061 16,035 9,316 Athletes (FE) 12,181 8,408 5,856 Events (FE) 438 523 531 Clusters (FE) 178 178 176 RMSE 0.771 1.161 1.176 Adj. $R^{2}$ 0.641 -0.166 -0.674 Within $R^{2}$0.905 -5.925 -8.453 First Stage F-stat 1,115.6 51.4 12.7 First Stage $p$-value $<2.2\times 10^{-16}$ $8.23\times 10^{-13}$ $3.64\times 10^{-4}$ Wu-Hausman $p$-value $<2.2\times 10^{-16}$ $<2.2\times 10^{-16}$ $<2.2\times 10^{-16}$ Notes: Two-stage least squares (2SLS) estimates of $\log(r_{i}+1)$, where $r_{i}$ is the centered finishing rank of athlete $i$ within event. The endogenous regressor is the athlete’s drafter position $p_{i}$, instrumented using the cluster-level peer mean swim performance (leave-one-out). No additional covariates are included. Fixed effects for athlete, cluster, and event are used. Standard errors are clustered at the event level. In Column (1), the coefficient is small but statistically significant, indicating modest benefits from trailing farther behind. The magnitude and significance of the effect increase in Columns (2) and (3), which restrict to smaller drafting clusters. In these subsamples, being positioned farther back in the pack is associated with significantly better performance, consistent with drafting theory. The precision of the estimates improves in Column (2) due to stronger first-stage relevance (F = 51.4), though instrument strength is more limited in Column (3) (F = 12.7). All models pass the Wu-Hausman test, supporting the endogeneity of drafter position. A total of 52,222 and 30,113 observations were dropped in Columns (2) and (3), respectively, due to missing or invalid values of the dependent variable, including cases where $\log(r_{i}+1)$ is undefined (e.g., due to negative centering). Significance levels: ${}^{***}p<0.01$, ${}^{**}p<0.05$, ${}^{*}p<0.1$
A.2 Robustness: Leave-one-out Peer-IV 2SLS Regressions under different sample sizes and model specifications
@@
A.3 Robustness: Leave-one-out Peer-IV 2SLS Estimates Using Theory-Implied Drafting Benefit Function

### v2

@@ Abstract
This paper examines strategic effort and positioning choices resulting in bandwagon effects under externalities in finite multi-stage games using causal evidence from triathlon (Reichel , 2025). Focusing on open-water swim drafting—where athletes reduce drag most effectively by swimming directly behind peers—we estimate its performance effects through a structural contest framework with endogenous, deterministic effort and drafting position. Leveraging exogenous variation from COVID-19 drafting bans in Austrian triathlons, we apply a panel leave-one-out (LOO/LOTO) group ability instrumental variables (IV) strategy to isolate the causal non-linear effect of drafting. Results from restricted sample analysis and pooled estimated bandwagon IV effects show substantial and nonlinear gains: in small (group size $<$ 10) drafting swim groups/clusters, each deeper position improves finishing rank on average by over 30%, with rapidly diminishing returns in larger groups. Leading however is consistently more costly than optimal positioning, aligning with theoretical predictions of energy expenditure (metabolic costs).
@@ 1 Introduction
This paper investigates the performance implications of open water swim drafting using rich panel data from triathlon races by modeling drafting in the first segment as a two-stage deterministic game. In the first stage, athletes exert some deterministic swimming effort; in the second stage, they position themselves optimally within a group, potentially benefiting heterogeneously from reduced drag behind other swimmers due to the inversly decaying nature of hydrodynamic benefits and rare overtakings (resorting) during the swim segment (Reichel (2025)).
Empirical estimates of drafting effects in the swim segment are still limited by measurement challenges. Group formations are likely endogenous and cannot be observed directly, and drafting exposure may correlate with unobserved ability (sorting/selection bias). Laboratory and biomechanical studies (e.g., Chatard (1990, 2003); Delextrat (2003); Bolon (2020)) focus on energy expenditure (metabolic) and performance gains from drafting, but large-scale causal estimates from race data have only been documented recently (in Reichel (2025)).
In our theoretical framework, we extend contest theory and incorporate models of strategic externalities (Tullock (1980); Moldovanu (2001); Konrad (2009); Che (2003)) to develop a structural model in which athletes first select a deterministic level of swimming effort, followed by an endogenous choice of drafting position. Drafting benefits are captured through a non-linear function that exhibits diminishing marginal returns beyond a certain positional threshold. The structural implications of the model are then integrated into an empirical two-stage least squares (2SLS) estimation framework. To address endogeneity in within-group drafting positions, we employ a leave-one-out (LOO/LOTO) average of peer swimming ability as an instrumental variable (IV).
@@ 2.2 Drafting Group Construction
Drafting groups can not be observed directly in the data. We thus infer drafting clusters using a single-linkage hierarchical agglomerative clustering algorithm approach applied to swim-out segment times within each event. Athletes are grouped into the same swim cluster if their exit times fall within a 5-second distance thresholds recorded. Earlier trials with complete linkage and wider thresholds (10–15 seconds) produced overly broad or unstable groupings. (Reichel (2025))
@@ 3.1 A Two-Stage Finite Game
This structure reflects a dynamic contest framework with endogenous positioning and group-related externalities (see Tullock (1980); Konrad (2009); Moldovanu (2001); Tadelis (2013); Vives (2005)).
@@ 3.1.1 Effort Cost Function
Assumption 1.
For the sake of simplicity we assume a linear cost function of effort:
@@ 3.2.2 Second Derivative and Convexity
Proposition 1 (Interior Optimal Drafting Position).
There exists a unique $d_{i}^{*}>3$ such that disutility is minimized, i.e., drafting benefit is maximized net of performance shortfall risk.
Proof.
  • For $d_{i}\leq 3$, $B(d_{i})=0\Rightarrow$ disutility is flat.
  • For $d_{i}>3$, $DIS_{i}$ is strictly convex.
  • Since $DIS_{i}\to\alpha c\bar{e}_{i}^{1}-(1-\alpha)\gamma$ as $d_{i}\to\infty$, the marginal gains vanish.
Hence, the minimum is interior and finite. ∎
@@ 3.2.3 Empirical Mapping
or to estimate effects. We can address address endogeneity by instrumenting $d_{i}$ using a leave-one-out group ability IV.
@@ 3.3 Why the Optimal Position Is Not the Last
Corollary 1.
Athletes face trade-offs — drafting increases with depth, but so do tactical disadvantages. Therefore, the optimal position is interior — balancing maximum benefit with control. This matches observed behavior in collective movement and contests with crowding externalities Hillman (1989); Krueger (1974).
@@ 4.1 Research Objective
Because drafting positions reflect strategic sorting and group formation dynamics, naive OLS is likely biased. We thus use a two-stage least squares (2SLS) framework with a leave-one-out group swim-time instrument to estimate a local average treatment effect (LATE) of drafting benefits (functional form from the theoretic model) or drafting positions on performance.
@@ 4.3 Endogeneity and Identification Strategy
We address this using a constructed leave-one-out group swim-time instrument $Z_{ie}$ to generate the relevant and exogenous variation in swim group drafting structure.
@@ 4.3.1 Leave-One-Out (LOO/LOTO) Instrumental Variable (IV)
Definition 1 (LOO/LOTO Instrument).
We define the instrument as:
@@
This captures the swim ability of groups in athlete $i$’s drafting within the current group, excluding $i$. This leave-one-out (LOO/LOTO) instrument captures the average group ability while avoiding mechanical reflection bias. By excluding athlete $i$, we remove the mechanical correlation between the instrument and the outcome. This strategy aligns with methods used in social interaction and group effect models to address endogeneity, e.g., Bramoulle (2009); Angrist (2014). Related applications include De Loecker (2011), who employ LOO instruments in production function estimation, and Guryan (2008), who instrument for group characteristics using group averages excluding the individual.
@@ 4.3.2 Drafting Benefit Transformation from Theoretic Model
4.4 Empirical Specification: Two-Stage Least Squares (2SLS) group-IV Estimation
To identify the causal effect of within-group drafting benefit on race performance, we estimate a two-stage least squares (2SLS) model using a leave-one-out group ability instrument. This setup accounts for endogeneity in observed drafting positions, which are determined by both athlete ability and strategic group formation.
@@ 4.4 Empirical Specification: Two-Stage Least Squares (2SLS) group-IV Estimation
We begin by estimating the drafting position as a function of group swim ability and controls:
@@
  • $Z_{ie}$: leave-one-out group average swim time (excluding athlete $i$);
@@
Using the fitted drafting benefit $\widehat{B}_{ieg}=f(\widehat{D}_{ieg})$ to validate the theoretical model or using $\widehat{D}_{ieg}$ directly, we estimate:
@@
4.4.1 Identification Strategy and Exogeneity
This 2SLS framework identifies the local average treatment effect (LATE) of drafting benefit for compliers—athletes whose position is affected by group group composition.
Assumption 2 (Identification Assumptions).
@@ 4.4.1 Identification Strategy and Exogeneity
  • Instrument exogeneity: $\mathbb{E}[\varepsilon_{ieg}\mid Z_{ie},S_{ie},\bm{X}_{ie},\lambda_{e},\alpha_{i},\alpha_{g}]=0$;
  • Monotonicity: An increase in group ability (faster groups) weakly shifts athletes to deeper drafting positions.
@@ 4.4.2 Structural Validation and Interpretation
The coefficient $\beta_{1}$ reflects the average marginal return to drafting benefit. The increasing and concave structure implies that:
  • Greatest positive net gains occur when transitioning from front positions backwards;
  • Diminishing returns occur deeper into the pack;
  • The optimal position is interior, as predicted by the theoretical model.
Empirical evidence of this shape supports the game’s interior equilibrium logic and nonlinear (dis-)utility (drafting benefit structure) topology.
@@ 4.5 Potential Extension: Projected Panel IV
For robustness and within-athlete variation, we define a projected instrument using historical group characteristics from the largest group containing athlete $i$, e.g.:
@@ 5 Empirical Results
Proposition 2.
In Table 3, the coefficient on drafting position is $-0.011$ ($p<0.1$) in the full sample (without restricting the outcome variable), implying that moving one position deeper in the drafting line improves expected rank by approximately 1.1%:
@@
This effect increases substantially in smaller groups. In drafting clusters with fewer than 10 athletes, the coefficient becomes $-0.972$ ($p<0.01$), translating to a 62.1% improvement in finishing rank per position:
@@
In very small clusters (group size $<5$), the coefficient is $-2.378$ ($p<0.1$), corresponding to a 90.7% improvement:
@@ 5.1 Using the theory-based functional form
with some uncalibrated $\lambda=0.5$ for the simple sake of simplicity and brevity. This specification reflects increasing but diminishing returns to deeper drafting positions.
Observation 1.
Across all models, the coefficient on $B(d_{i})$ is significantly negative, consistent with improved performance from stronger drafting effects.
The effect is strongest in smaller groups: in groups with fewer than 10 athletes, a one-unit increase in drafting benefit improves log-rank by approximately 7.1 to 9.6 points (Columns 5–6). In broader samples, the benefit remains significant and negative—$-6.65$ in the full sample (Column 1), and $-4.84$ to $-3.61$ in groups smaller than 20 (Columns 3–4).
@@ 6 Conclusion
This paper contributes comprehensive analysis on Strategic Effort and Bandwagon Effects in Finite Multi-Stage Games with Non-linear Externalities and draws on empirical causal evidence using rich panel data from drafting in endurance sports competitions, predicting effects on final stage rank outcomes focusing on competitive athletes. We develop and estimate log-linear models with endogenous drafting position, using a leave-one-out group swim ability as an instrument (Angrist , 2009; Manski , 1993; degiorgi2020group). The results validate the strategic and nonlinear nature of drafting benefits in multi-stage competitions such as triathlon.
@@ Appendix A Appendix
A.1 Leave-one-out Group-IV 2SLS Regressions with Capped/Limited Outcome Variable Rank $<$ 250
@@
(1) Full Sample (2) Group Size $<$ 10 (3) Group Size $<$ 5 Drafter Position (Fit) -0.011 -0.972∗∗∗ -2.378 (0.0053) (0.185) (1.073) Observations 34,061 16,035 9,316 Athletes (FE) 12,181 8,408 5,856 Events (FE) 438 523 531 Clusters (FE) 178 178 176 RMSE 0.771 1.161 1.176 Adj. $R^{2}$ 0.641 -0.166 -0.674 Within $R^{2}$0.905 -5.925 -8.453 First Stage F-stat 1,115.6 51.4 12.7 First Stage $p$-value $<2.2\times 10^{-16}$ $8.23\times 10^{-13}$ $3.64\times 10^{-4}$ Wu-Hausman $p$-value $<2.2\times 10^{-16}$ $<2.2\times 10^{-16}$ $<2.2\times 10^{-16}$ Notes: Two-stage least squares (2SLS) estimates of $\log(r_{i}+1)$, where $r_{i}$ is the centered finishing rank of athlete $i$ within event. The endogenous regressor is the athlete’s drafter position $p_{i}$, instrumented using the cluster-level Group mean swim performance (leave-one-out). No additional covariates are included. Fixed effects for athlete, cluster, and event are used. Standard errors are clustered at the event level. In Column (1), the coefficient is small but statistically significant, indicating modest benefits from trailing farther behind. The magnitude and significance of the effect increase in Columns (2) and (3), which restrict to smaller drafting clusters. In these subsamples, being positioned farther back in the pack is associated with significantly better performance, consistent with drafting theory. The precision of the estimates improves in Column (2) due to stronger first-stage relevance (F = 51.4), though instrument strength is more limited in Column (3) (F = 12.7). All models pass the Wu-Hausman test, supporting the endogeneity of drafter position. A total of 52,222 and 30,113 observations were dropped in Columns (2) and (3), respectively, due to missing or invalid values of the dependent variable, including cases where $\log(r_{i}+1)$ is undefined (e.g., due to negative centering). Significance levels: ${}^{***}p<0.01$, ${}^{**}p<0.05$, ${}^{*}p<0.1$
A.2 Robustness: Leave-one-out Group-IV 2SLS Regressions under different sample sizes and model specifications
@@
A.3 Robustness: Leave-one-out Group-IV 2SLS Estimates Using Theory-Implied Drafting Benefit Function